### v1

# Generalizable Deepfake Detection via Effective Local-Global Feature Extraction

Abstract
The rapid advancement of GANs and diffusion models has led to the generation of increasingly realistic fake images, posing significant hidden dangers and threats to society. Consequently, deepfake detection has become a pressing issue in today’s world. While some existing methods focus on forgery features from either a local or global perspective, they often overlook the complementary nature of these features. Other approaches attempt to incorporate both local and global features but rely on simplistic strategies, such as cropping, which fail to capture the intricate relationships between local features. To address these limitations, we propose a novel method that effectively combines local spatial-frequency domain features with global frequency domain information, capturing detailed and holistic forgery traces. Specifically, our method uses Discrete Wavelet Transform (DWT) and sliding windows to tile forged features and leverages attention mechanisms to extract local spatial-frequency domain information. Simultaneously, the phase component of the Fast Fourier Transform (FFT) is integrated with attention mechanisms to extract global frequency domain information, complementing the local features and ensuring the integrity of forgery detection. Comprehensive evaluations on open-world datasets generated by 34 distinct generative models demonstrate a significant improvement of 2.9% over existing state-of-the-art methods.
1 Introduction
With the rapid advancement of deep learning, generative models such as GANs and diffusion models have revolutionized content creation, finding widespread applications across industries including entertainment and media. However, this progress has also introduced significant challenges. The increasing sophistication of deepfake technology has allowed the generation of highly realistic fake images and videos, making them increasingly difficult to detect and posing serious risks to societal trust, individual privacy, and security.
In response, various deepfake detection mechanisms have been conceived. Early methods primarily focused on analyzing image features, such as texture, color, and lighting Yu et al. (2020); Liu et al. (2020), to identify forgery traces. However, these approaches were often limited to training and testing within a single domain, overlooking the critical issue of cross-domain generalization. In real-world scenarios, it is essential for detection models to effectively generalize to unseen deepfake sources, including those involving unfamiliar generative models or novel categories.
Figure 1: Comparison with Previous Methods. Previous methods often extract only local information or global information to detect artifact information. However, these methods are not enough to fully capture effective artifact information. Recent methods also combine local and global information to detect forged images, but they only extract local information through simple cropping, destroying the relationship between local information. Compared with them, our method uses DWT and sliding window tiling to reasonably and effectively extract local time-frequency information, and combines the phase part of FFT to extract global frequency domain information.
In recent years, substantial efforts have been devoted to forgery detection, particularly to identifying deepfake sources that are challenging to detect in real-world scenarios, as illustrated in Figure 1 (a). To enhance generalization performance, some studies emphasize fine-grained features, such as local artifacts Haliassos et al. (2021); Chen et al. (2022) and the relationships between pixel points Chen et al. (2021); Tan et al. (2024b). Others focus on global artifact traces, including global spatial information Tan et al. (2023); Ojha et al. (2023) and frequency-domain features Tan et al. (2024a). However, prioritizing either local or global information in isolation may result in the loss of complementary details, thereby limiting generalization capabilities. To address this limitation, some methods integrate both local and global features. For example, Chen and Yang (2020) extract local features through clipping in the spatial domain while leveraging global features in the frequency space. Similarly, Wu et al. (2024) employ a local attention module and facilitates interaction with features from distant adjacent PPG graphs to capture local and global information, respectively. However, these approaches often rely on coarse image block segmentation for local feature extraction, which can lead to forced separation and loss of relevant feature relationships.
Accordingly, we propose a simple yet efficient feature extractor to capture global frequency domain and local spatial-frequency domain information, as illustrated in Figure 1 (b), which is plug-and-play and can be integrated with any CNN classifier. Specifically, our method introduces two key modules, DWT-Based Local Spatial-Frequency Domain Feature Capture Block (LoSFB) and FFT-Based Global Frequency Domain Feature Capture Block (GloFB), both using the Window Attention Block as the feature extractor. After performing the DWT, we concatenate features from four different frequency bands, each with dimensions $h\times w$. Then a $2\times 2$ sliding window is applied to extract the tiled features from each frequency band. Finally, the features within the window are tiled again, resulting in new features with dimensions $4\times(h\times w)$. At this time, when we extract the $4\times 4$ window features, they not only obtain features in four frequency domains, but also retain the local spatial-frequency information within the window, greatly improving the effect of artifact detection. Additionally, we incorporate the phase information from the FFT as a complement to the DWT, enabling the extraction of global frequency domain information. This helps guide the network to effectively learn global artifact patterns and identify more comprehensive traces of forgery. To fully assess the generalizability of our method, we evaluate it on large image datasets generated by 34 distinct models [^1] ProGAN, StyleGAN, StyleGAN2, BigGAN, CycleGAN, StarGAN, GauGAN, Deepfake, WFIR, SAN, AttGAN, BEGAN, CramerGAN, InfoMaxGAN, MMDGAN, RelGAN, S3GAN, SNGAN, STGAN, ADM, DDPM, IDDPM, LDM, PNDM, VQDiffusion, Stable-Diffusion v1, Stable-Diffusion v2, DALLE, Glide, Midjourney, Stable-Diffusion v1.4, Stable-Diffusion v1.5, Wukong, DALLE-2. Comprehensive experiments demonstrate that our method generalizes well across diverse and unseen sources, demonstrating a significant improvement over state-of-the-art methods by 2.9%. Our paper makes the following contributions:
  • We propose a plug-and-play feature extractor that seamlessly integrates local spatial-frequency and global frequency-domain information. This method can be effortlessly incorporated into any CNN classifier and effectively improves the model’s generalization capability.
  • We use sliding window interpretability to extract local spatial frequency information, and creatively use only phase spectrum to obtain global features to capture broader forgery artifacts.
  • Our experiments validate the effectiveness of the proposed method, which exhibits strong generalization across image detection tasks involving 34 different generation models. The method achieves a notable 2.9% improvement over state-of-the-art approaches, highlighting its superiority and robustness.
2 Related Work
In this section, we present a concise review of deepfake detection methods, mainly dividing them into three main groups: deepfake detection based on local information, deepfake detection based on global information, and deepfake detection based on local and global information fusion.
2.1 Deepfake Detection Based on Local Information
Early researches on deepfake detection often focus on specific facial regions, such as the eyes and lips, to identify forgery traces Haliassos et al. (2021); Chen et al. (2022). However, as forgery techniques have advanced, these simple traces have been increasingly refined and obscured, rendering biometric-based detection methods insufficient for identifying sophisticated forged images. Experiments have shown that forgery traces are often embedded in fine-grained details, prompting a shift in focus toward extracting local artifact information. For instance, Chai et al. (2020) use limited receptive fields to capture local features, while Zhao et al. (2021) employ multiple spatial attention heads to enhance the network’s focus on distinct local regions. Other studies analyze forgery traces by examining relationships between local pixels. Chen et al. (2021) propose learning the similarity of local regions for a fine-grained assessment of patch similarity across scales, and Tan et al. (2024b) improve forgery detection generalization by identifying upsampling artifacts in nearby pixel grids. While capturing local artifact details, these methods overlook the fact that the overall image may still contain forged features, which are essential for generalization.
2.2 Deepfake Detection Based on Global Information
Early studies leveraged spatial information from generated images to detect fake content. For example, Yu et al. (2020); Liu et al. (2020) distinguish real and fake images using global information, such as color space and global texture. Some studies Yu et al. (2019) use GANs to generate global fingerprint information that may assist in detection. Additionally, Tan et al. (2023) and Ojha et al. (2023) extract artifact representations using global feature maps and global feature gradients, respectively. Moreover, since GAN-generated images often rely on upsampling operations that introduce artifacts at the frequency domain level, several studies have explored forgery detection in the frequency domain. For instance, Tan et al. (2024a) use FFT to extract global high-frequency information, uncovering more refined forgery features. While these methods capture sufficient global artifact information, they overlook the fact that effective artifact features often reside in subtle, localized regions. The neglect of local artifact information can significantly impair the generalization capabilities of forgery detection.
2.3 Deepfake Detection Based on Local and Global Information Fusion
Whether relying on global or local information alone, these methods detect forgery using only a single category of features, overlooking the significant information gain that arises from combining complementary features at different levels. Chen and Yang (2020) capture local features through spatial domain cropping and global features in the frequency domain. Xue et al. (2023) employ separate branches for local area feature extraction and global feature extraction, using the latter to obtain residual features. Wu et al. (2024) introduce a local attention module that interacts with distant neighboring PPG graphs to gather both local and global information. While these studies incorporate both global and local information, their approach to processing local information is limited by artificial segmentation of the face into regions, which leads to the loss of important forged details.
3 Main Backbone
The overall structure of our method is shown in Figure 2. Inspired by Mao et al. (2024), we propose an efficient local-global feature extraction framework designed to capture both global and local information. Specifically, the framework consists of two main branches: LoSFB and GloFB. Both branches utilize the Window Attention Block to extract the most relevant features, which are then fused and passed to the classifier for final classification. Notably, our approach is compatible with any classifier, allowing us to focus primarily on the feature extraction process.
Figure 2: Architecture of Our Method for Generalizable Deepfake Detection. In order to capture more comprehensive and reasonable forgery features, we use LoSFB and GloFB to capture local and global information, respectively. Specifically, we use DWT and sliding window tiling to construct local spatial-frequency features, and only use the phase part of FFT to construct global frequency domain features. Both of them use Window Attention Block to extract features. In addition, in order to weaken the information weight of global features, we add hyperparameters $\lambda$ to the global features when merging the two types of features. The final extracted features are passed into a classifier.
3.1 Window Attention Block
The Window Attention Block (WAttention) consists of two main components: (i) the Local Attention Block (LoAttention) and (ii) the Global Window MLP Layer (GMLP). Before feeding the features into LoAttention and GMLP, we preprocess them to enhance their representational quality. A $1\times 1$ convolution layer is applied to extract cross-channel features, following by a $3\times 3$ depthwise separable convolution to better capture the local features within each channel. The resulting convolved features are denoted as $X_{in}\in{\mathbb{R}}^{B\times C\times H\times W}$.
LoAttention. Unlike traditional attention mechanisms, which are typically designed to capture global information, we employ a sliding window approach that divides the features into windows of size $b\times b$. By restricting the attention module to these local windows, the focus shifts to effectively extracting local features. The features are then divided into $Q,K,V\in{\mathbb{R}}^{S\times C\times N}$, where $S=B\times\frac{H}{b}\times\frac{W}{b}$ and $N=b\times b$, as required by the attention module. Since the attention operates locally within each window, the features within window $i$ are represented as $Q_{i},K_{i},V_{i}\in{\mathbb{R}}^{C\times N}$. These localized features are fed into the attention module for computation:
$A_{i}=\text{Attention}(Q_{i},K_{i},V_{i})=\text{SoftMax}(Q_{i}K_{i}^{\top})V_{i}.$ (1)
The results of all window attention operations can be expressed as $A=\{A_{1},A_{2},\cdots,A_{S}\}\in{\mathbb{R}}^{S\times C\times N}$.
GMLP. To preserve global correlations while capturing fine-grained features, we integrate a global MLP within each window. This aims to address the limitations of local windows in modeling global relationships, ensuring a seamless fusion of local details and global context. Specifically, we reuse the value $V_{i}$ from LoAttention and process it through the global MLP, applying a nonlinear transformation to enhance its representational capacity. The result $\hat{V_{i}}$ is formally expressed as:
$\text{GMLP}(V_{i})=G(\text{Linear}(V_{i})),$ (2)
where $G(\cdot)$ denotes the GELU function. The GMLP recombines and redistributes the local window information, enabling the mixing of local frequency domain features and enhancing the learning of global features.
After combining the outputs from LoAttention and GMLP via a dot product, we use the inverse window operation to restore the feature to its original size $B\times C\times H\times W$, followed by a convolution to map the features back to the original feature dimensions. The final output of the WAttention can be expressed as:
$\text{X}_{W}=\text{conv}(\sum_{i=1}^{s}\text{Attention}(Q_{i},K_{i},V_{i})\cdot\sum_{i=1}^{s}G(\text{Linear}(V_{i}))).$ (3)
3.2 DWT-Based Local Spatial-Frequency Domain Feature Capture Block
Previous methods often leverage the Fast Fourier Transform (FFT) or Discrete Cosine Transform (DCT) to capture relationships between adjacent nodes in the frequency domain. However, these approaches can lead to the loss of positional information, which is crucial for extracting detailed features. To overcome this limitation, we utilize the Discrete Wavelet Transform (DWT), which decomposes the image into the spatial-frequency domain, represented as:
$\text{DWT}(I)=\{I_{LL},I_{LH},I_{HL},I_{HH}\},$ (4)
where $I_{LL},I_{LH},I_{HL},I_{HH}\in{\mathbb{R}^{C\times\frac{H}{2}\times\frac{W}{2}}}$ represent the low-frequency component of the input and the high-frequency components in the vertical, horizontal, and diagonal directions, respectively. To obtain more comprehensive information, we fuse the data from the four frequency domains, resulting in new input features $\hat{I}\in{\mathbb{R}^{4\times C\times\frac{H}{2}\times\frac{W}{2}}}$. As illustrated in Figure 3, to extract detailed features, we divide each channel’s features into $2\times 2$ windows, with each window capturing localized spatial-frequency information within its respective frequency band. Each window is treated as a small square and tiled into rows within each band. This arrangement ensures that features within the $2\times 2$ windows are preserved in a structured format. Then we tile the features in the window here by rows. At this time, the features from each frequency band are tiled into one row in a specific manner, such that the four adjacent feature values correspond to the $2\times 2$ window from the original image. Additionally, as shown in Figure 3, we integrate features from all frequency bands to generate new input features ${\hat{I}}_{\text{input}}\in{\mathbb{R}}^{C\times 4\times\frac{H\times W}{4}}$. To further enhance the network’s performance, we apply LayerNorm Lei Ba et al. (2016), which ensures that the input data remains within a stable range, reduces dependency on initialization, and improves the stability of model training.
The feature ${\hat{I}}_{\text{input}}$ is then fed into WAttention with a sliding window size set to $b=4$. As depicted in Figure 3, within each $4\times 4$ window, the representation includes not only the four adjacent feature values from the original image but also the feature information across four distinct frequency bands. This configuration provides WAttention with comprehensive and fine-grained spatial-frequency domain information, enabling more detailed feature extraction. Following this process, we apply the IDWT to reconstruct the features back into spatial domain representations, denoted as ${\hat{I}_{\text{DWT\_output}}}$.
3.3 FFT-Based Global Frequency Domain Feature Capture Block
Although artifact information is often concentrated in finer details, forged features can also be embedded in the global information of an image, remaining imperceptible to the naked eye. While DWT and sliding window tiling are effective for extracting local features, we also use FFT capture global frequency domain information from a complementary perspective. In this module, FFT is employed to transform the image into the frequency domain at the pixel level, enabling the extraction of global frequency domain features. Inspired by Zhao et al. (2024), we recognize that amplitude primarily reflects properties such as color and brightness, while phase encodes finer details, including potential artifacts. Based on this insight, we focus on extracting the phase features as input to WAttention, setting the window size to $b=8$. The amplitude remains unmodified during this process. After training the phase features, we use IFFT to recombine them with the original amplitude, transforming the features back into the spatial domain. The resulting features are denoted as ${\hat{I}_{\text{FFT\_output}}}$.
Finally, we set a hyperparameter $\lambda$ to minimize the impact of global information extraction. The final feature is:
${\hat{X}_{\text{output}}}=(1-\lambda)*\hat{I}_{\text{DWT\_output}}+\lambda*\hat{I}_{\text{FFT\_output}}.$ (5)
Figure 3: After DWT, sliding window tiling is used to tile the features of each channel into a shape of $4\times(h\times w)$. In the subsequent LoAttention module, $4\times 4$ features are extracted as shown in the blue box, which not only contains local spatial frequency information, but also contains all different frequency band features.
4 Experiments
To verify the excellent deepfake detection performance of our method, we followe the paradigms of Ojha et al. (2023) and Wang et al. (2020).
4.1 Dataset
Training set. To ensure a consistent basis for comparison, we use the training set of ForenSynthsWang et al. (2020). The training set consists of 20 different categories, each containing 18,000 synthetic images generated using ProGAN, and an equal number of real images from the LSUN dataset. Following the paradigms of Ojha et al. (2023), we adopt specific 4-class training settings, denoted as (car, cat, chair, horse).
Testing set. In order to comprehensively compare the effectiveness of our proposed method, we select different GAN and diffusion datasets for testing, which contain a wide range of real-world images.
  • ForenSynths: The test set comprises 10 subsets, including: ProGAN Karras (2017), StyleGAN Karras (2019), StyleGAN2 Karras et al. (2020), BigGAN Brock (2018), CycleGAN Zhu et al. (2017), StarGAN Choi et al. (2018), GauGAN Park et al. (2019), Deepfake Rossler et al. (2019), WFIR and SAN Dai et al. (2019).
  • GANGen-Detection Tan et al. (2024a): To evaluate more realistic scenarios, we extend our evaluation using 9 additional GAN-generated images. There are 4K test images for each model, with equal numbers of real and fake images.
  • DiffusionForensics Wang et al. (2023): To expand the testing scope, we adopt the diffusions dataset of DIRE Wang et al. (2023) for evaluation, including ADM Dhariwal and Nichol (2021), DDPM Ho et al. (2020), IDDPM Nichol and Dhariwal (2021), LDM Rombach et al. (2022), PNDM Liu et al. (2022), VQ-Diffusion Gu et al. (2022), Stable Diffusion v1 Rombach et al. (2022), Stable Diffusion v2 Rombach et al. (2022). The real images are sampled from LSUN Yu et al. (2015) and ImageNet Russakovsky et al. (2015) datasets.
  • UniversalFakeDetect Ojha et al. (2023): This test set contains images generated from Glide Nichol et al. (2021), DALL-Emini Ramesh et al. (2021), LDM Rombach et al. (2022). It adopts images of LAION Schuhmann et al. (2021) and ImageNetRussakovsky et al. (2015) datasets as the real data.
  • AIGCDetectBenchmark Zhu et al. (2024): This dataset mainly adopts diffusion model to generate images, and is generated using more realistic advanced methods, including Midjourney, SDv1.4, SDv1.5 Rombach et al. (2022), ADM Dhariwal and Nichol (2021), GLIDE Nichol et al. (2021), Wukong, VQDM Gu et al. (2022), which is a huge challenge for deepfake detection.
4.2 Implementation Details
We design a lightweight CNN network using convolutional layers and Resnet network as classifiers. The network is trained using the Adam optimizerKingma (2014) with a learning rate of $2\times 10^{-4}$. The batchsize is set to 128, and we train the network for 90 epochs. We also use a learning rate decay strategy that reduces the learning rate by twenty percent after every ten epochs. What’s more, the hyperparameter $\lambda$ is set to 0.2. Our approach is implemented using PyTorch on an NVIDIA RTX A6000 GPU. To evaluate the performance of the proposed method, we follow the evaluation metrics used in baselines Ojha et al. (2023); Tan et al. (2024a), which include average precision score (A.P.) and accuracy (Acc.).
4.3 Quantitative analysis
We compare with the previous methods: CNNDetection Wang et al. (2020), Patchfor Chai et al. (2020), F3Net Qian et al. (2020), FrePGAN Jeong et al. (2022), GANDetection Mandelli et al. (2022), LGrad Tan et al. (2023), Ojha Ojha et al. (2023), FreqNet Tan et al. (2024a) and NPR Tan et al. (2024b). We conduct two and three sets of experiments on GAN and diffusion respectively.
GAN. As shown in Table 1 and 2, our detection on GAN has achieved good results. Table 1 presents a detailed comparison of detector performance on the test set of ForenSynths Wang et al. (2020), highlighting higher mean accuracy (Acc.) and comparable mean average precision (A.P.) metrics. For unavailable test results, we use ”-” as a placeholder and exclude these values from the average calculations. Our method achieves an average accuracy of 86.9%, which is 1% and 1.7% higher than the latest methods, NPR and FreqNet, underscoring its effectiveness in generalizable deepfake detection. Furthermore, to evaluate cross-GAN-source generalization, we conducted tests on the GANGen-Detection dataset Tan et al. (2024a), as summarized in Table 2. Our method achieves an impressive average accuracy of 96.5%, significantly outperforming the best-performing baselines, NPR Tan et al. (2024b) and FreqNet Tan et al. (2024a), which achieved 94.8% and 94.0%, respectively.
Diffusion. To further evaluate the generalization ability of our method, we train on ProGAN and test on images generated by various diffusion models. The results are presented in Tables 3 and 4. Despite being trained exclusively on ProGAN-generated images, our method demonstrates strong generalization capabilities across diverse diffusion models. Specifically, it achieves mean Accuracy (Acc.) and mean Average Precision (A.P.) of 98.5% and 99.9%, respectively. Compared to NPR and FreqNet, our method improves mean Acc. by 2.9% and 17.8%, respectively. Additionally, on the UniversalFakeDetect dataset Ojha et al. (2023), our method achieves an average accuracy of 95.0%, representing a notable improvement over existing approaches.
With the rapid advancements in tools like Midjourney and diffusion models, increasingly realistic fake images are being generated, posing significant challenges for deepfake detection. To address this, we conduct additional evaluations on the AIGCDetectBenchmark dataset Zhu et al. (2024), excluding some subsets of data previously tested. The results, summarized in Table 5, highlight the robustness of our approach. Our method achieves a notable improvement in mean accuracy, outperforming NPR and FreqNet by 3.7% and 7.8%, respectively, and reaching an accuracy of 77.6%.
Method ProGAN GauGAN StyleGAN2 BigGAN StyleGAN CycleGAN StarGAN Deepfake WFIR SAN Mean
Acc. A.P. Acc. A.P. Acc. A.P. Acc. A.P. Acc. A.P. Acc. A.P. Acc. A.P. Acc. A.P. Acc. A.P. Acc. A.P. Acc. A.P.
CNNDetection 91.4 99.4 63.9 92.2 76.4 97.5 52.9 73.3 63.8 91.4 72.7 88.6 63.8 92.2 51.7 62.3 91.7 96.3 48.7 59.5 67.7 85.3
Patchfor 97.8 100.0 57.2 55.4 83.6 98.5 64.7 69.5 82.6 93.1 74.5 87.2 100.0 100.0 85.0 93.2 - - 75.3 76.3 80.1 85.9
F3Net 99.4 100.0 58.1 56.1 88.0 99.8 65.3 69.9 92.6 99.7 76.4 84.3 100.0 100.0 63.5 78.8 - - 47.3 46.7 76.7 81.7
FrePGAN 99.0 99.9 60.3 77.1 84.1 98.6 69.2 71.7 80.7 89.6 71.1 74.4 99.9 100.0 70.9 91.9 - - - - 79.3 87.9
LGrad 99.9 100.0 72.4 79.3 96.0 99.9 82.9 90.7 94.8 99.9 85.3 94.0 99.6 100.0 58.0 67.9 60.5 65.1 50.0 51.4 79.9 84.8
Ojha 100.0 100.0 99.7 100.0 75.7 97.8 95.1 99.3 84.4 97.1 98.7 99.8 99.9 99.4 67.4 82.0 64.2 68.7 56.5 82.2 84.2 92.6
FreqNet 99.6 100.0 93.4 98.6 88.0 99.5 90.5 96.0 90.2 99.7 95.8 99.6 85.7 99.8 88.9 94.4 48.0 49.6 71.9 80.1 85.2 91.4
NPR 99.9 100.0 80.9 83.0 99.6 100.0 84.0 85.6 98.1 99.8 95.2 98.1 99.8 100.0 77.2 76.0 60.6 62.9 64.4 66.0 85.9 87.1
ours 99.9 100.0 86.3 97.4 99.9 100.0 87.1 93.6 99.9 100.0 97.1 98.9 90.8 95.9 64.5 72.2 77.8 87.9 65.6 70.6 86.9 91.7
Table 1: Cross-GAN-Sources Evaluation on the test set of ForenSynths Wang et al. (2020). Bold and underline represent the best and second-best performance, respectively.
Method BEGAN SNGAN CramerGAN MMDGAN RelGAN STGAN S3GAN AttGAN InfoMaxGAN Mean
Acc. A.P. Acc. A.P. Acc. A.P. Acc. A.P. Acc. A.P. Acc. A.P. Acc. A.P. Acc. A.P. Acc. A.P. Acc. A.P.
CNNDetection 50.2 44.9 62.7 90.4 81.5 97.5 72.9 94.4 53.3 82.1 63.0 92.7 55.2 66.1 51.1 83.7 71.1 94.7 62.3 82.9
Patchfor 97.1 100.0 97.6 99.8 97.8 99.9 97.9 100.0 99.6 100.0 92.7 99.8 66.8 68.8 68.0 92.9 93.6 98.6 90.1 95.4
F3Net 87.1 97.5 51.6 93.6 89.5 99.8 73.7 99.6 98.8 100.0 60.3 99.9 65.4 70.0 85.2 94.8 67.1 83.1 75.4 93.1
GANDetection 67.9 100.0 66.7 90.6 67.8 99.7 67.7 99.3 60.9 86.2 69.6 97.2 69.6 83.5 57.4 75.1 67.6 92.4 66.1 91.6
LGrad 69.9 89.2 78.0 87.4 50.3 50.3 57.5 67.3 89.1 99.1 54.8 68.0 78.5 86.0 68.6 93.8 71.1 82.0 68.6 80.8
Ojha 88.9 96.3 87.7 96.5 89.8 99.1 89.7 99.0 94.0 98.1 82.6 91.5 94.7 98.9 90.0 96.8 87.4 96.5 89.4 97.0
FreqNet 98.8 100.0 85.4 90.4 95.2 98.2 95.2 98.2 100.0 100.0 98.8 100.0 88.3 94.3 89.8 98.8 94.0 97.5 94.0 97.5
NPR 99.6 99.9 93.3 97.3 98.5 98.3 98.5 98.3 99.8 100.0 99.6 100.0 79.8 78.9 92.5 98.6 91.5 97.2 94.8 96.5
ours 97.7 99.8 96.4 99.6 98.2 99.8 98.8 99.9 98.8 88.8 99.4 100.0 83.4 90.2 98.7 99.8 96.8 99.7 96.5 98.7
Table 2: Cross-GAN-Sources Evaluation on the 9 GANs dataset Tan et al. (2024a).
Method DDPM VQ-Diffusion SDv1 ADM SDv2 IDDPM PNDM LDM Mean
Acc. A.P. Acc. A.P. Acc. A.P. Acc. A.P. Acc. A.P. Acc. A.P. Acc. A.P. Acc. A.P. Acc. A.P.
CNNDetection 62.7 76.6 50.0 71.0 38.0 76.7 53.9 71.8 52.0 90.3 50.2 82.7 50.8 90.3 50.4 78.7 51.0 90.3
Patchfor 62.3 97.1 100.0 100.0 90.7 99.8 77.5 93.9 94.8 100.0 50.0 91.6 50.2 99.9 99.5 100.0 78.1 97.8
F3Net 84.7 99.4 100.0 100.0 73.4 97.2 80.9 96.9 99.8 100.0 74.7 98.9 72.8 99.5 100.0 100.0 85.8 99.0
GANDetection 62.3 46.4 51.1 51.2 39.8 65.6 51.1 53.1 50.1 36.9 50.2 63.0 50.6 79.0 51.6 48.1 50.8 55.4
LGrad 99.9 100.0 96.2 100.0 90.4 99.4 86.4 97.5 97.1 100.0 66.1 92.8 69.5 98.5 99.7 100.0 88.2 98.5
Ojha 72.9 78.8 77.7 99.2 59.0 93.0 88.7 98.6 53.3 87.6 73.4 97.1 86.2 99.2 50.7 87.9 70.2 92.5
FreqNet 91.4 99.8 100.0 100.0 63.9 98.1 67.2 91.3 81.8 98.4 59.0 97.3 85.2 99.8 98.9 100.0 80.9 98.1
NPR 99.4 100.0 100.0 100.0 97.3 99.8 88.7 98.6 96.4 100.0 87.5 98.3 97.5 100.0 100.0 100.0 95.8 99.6
ours 99.4 100.0 99.9 100.0 95.6 99.7 96.3 99.4 99.2 100.0 99.4 100.0 99.4 100.0 99.8 100.0 98.7 99.9
Table 3: Cross-Diffusion-Sources Evaluation on the test of DiffusionForensics Wang et al. (2023).
Method Guided Glide_50_27 Glide_100_10 Glide_100_27 LDM_100 LDM_200 LDM_200_cfg dalle Mean
Acc. A.P. Acc. A.P. Acc. A.P. Acc. A.P. Acc. A.P. Acc. A.P. Acc. A.P. Acc. A.P. Acc. A.P.
CNNDetection 54.9 66.6 54.2 76.0 53.3 72.9 53.0 71.3 51.9 63.7 52.0 64.5 51.6 63.1 51.8 61.3 52.8 67.4
Patchfor 74.2 81.4 84.9 98.8 87.3 99.7 82.8 99.1 95.8 99.8 95.6 99.9 94.0 99.8 79.8 99.1 86.8 97.2
F3Net 69.2 70.8 88.5 95.4 88.3 95.4 87.0 94.5 74.1 84.0 73.4 83.3 80.7 89.1 71.6 79.9 79.1 86.5
GANDetection 50.1 51.0 51.7 53.5 51.2 52.6 51.1 51.9 54.7 65.8 54.9 65.9 53.8 58.9 67.2 83.0 54.3 60.1
LGrad 71.8 76.0 91.6 96.0 90.9 95.5 88.5 94.2 95.7 99.3 94.7 99.1 95.1 99.0 89.9 97.8 89.8 94.6
Ojha 69.7 87.6 79.1 99.8 77.9 94.5 78.5 95.3 95.0 99.3 99.0 99.3 74.0 92.5 87.3 97.5 82.6 95.7
FreqNet 67.3 75.7 86.7 96.3 87.9 96.4 84.5 96.0 97.9 99.9 97.5 99.9 97.4 99.9 97.4 99.8 89.6 95.5
NPR 74.8 78.8 98.3 99.8 98.7 99.7 98.2 99.8 98.8 99.8 99.0 99.8 98.8 99.7 95.9 99.7 94.8 97.0
ours 82.4 87.1 98.3 99.7 97.6 99.8 97.6 99.6 99.4 100.0 99.3 100.0 99.0 99.9 86.0 95.6 95.0 97.7
Table 4: Cross-Diffusion-Sources Evaluation on the diffusion test set of Ojha Ojha et al. (2023).
Method ADM Glide Midjourney SDv1.4 SDv1.5 VQDM Wukong DALLE2 Mean
Acc. A.P. Acc. A.P. Acc. A.P. Acc. A.P. Acc. A.P. Acc. A.P. Acc. A.P. Acc. A.P. Acc. A.P.
LGrad 65.8 70.2 60.6 73.8 64.6 70.3 65.9 68.7 66.3 69.1 67.2 70.6 63.2 65.0 66.9 86.9 65.1 71.8
Ojha 67.6 87.2 61.7 85.5 56.7 74.5 63.1 87.2 62.8 86.3 84.9 96.8 71.2 90.9 49.5 60.2 73.5 83.6
FreqNet 83.3 91.5 81.7 88.9 69.9 79.0 64.3 74.3 64.9 75.5 81.7 89.6 57.7 67.0 55.1 54.6 69.8 77.6
NPR 70.6 75.1 77.2 83.2 76.6 82.6 76.8 82.8 78.0 83.8 75.2 77.7 74.5 78.2 61.9 71.6 73.9 79.4
ours 80.2 83.7 80.8 93.0 73.0 79.4 78.4 84.1 77.8 87.9 83.2 88.7 75.5 78.0 71.3 89.6 77.6 85.6
Table 5: Cross-Diffusion-Sources Evaluation on the GenImage Dataset Zhu et al. (2024).
Method
Mean Acc. of
16 sub-testsets
CNNDetectionWang et al. (2020) 70.8
LGradTan et al. (2023) 75.3
OjhaOjha et al. (2023) 78.4
FreqNetTan et al. (2024a) 78.1
NPRTan et al. (2024b) 81.8
ours 84.7
Table 6: The mean accuracy of all 16 generation models on GenImage Zhu et al. (2024) dataset.
DWT Window Tiling FFT WAttention mean Acc. mean A.P.
$\surd$ 67.2 70.8
$\surd$ $\surd$ 49.5 52.3
$\surd$ $\surd$ 82.3 87.6
$\surd$ $\surd$ $\surd$ 83.9 90.2
$\surd$ $\surd$ $\surd$ $\surd$ 84.7 90.9
Table 7: Ablation study on the GenImage Zhu et al. (2024).
Conclusion. To comprehensively evaluate the superiority of our method, we test it on the entire GenImage dataset Zhu et al. (2024) and calculated the average accuracy. The GenImage dataset offers extensive coverage, including content generated by GANs as well as high-definition images from diffusion models. As shown in Table 6, our method achieves optimal detection performance, with an accuracy of 84.7%, significantly surpassing the accuracy of NPR and FreqNet, which are 81.8% and 78.1%, respectively.
Figure 4: Logit Distributions of Extracted Forgery Features. We compare the state of the art NRP Tan et al. (2024b) and our method, both tuned with 4-class ProGAN Karras (2017) data. A total of four testing GANs and diffusion models are considered, including ProGAN Karras (2017), StyleGAN Karras (2019), IDDPM Nichol and Dhariwal (2021) and VQDM Gu et al. (2022), each randomly sampled 1k real and 1k fake images.
4.4 Qualitative Analysis
To further assess the generalization ability of our method, we visualize the logit distributions of NPR and our approach, as shown in Figure 4. This visualization highlights how effectively each trained model distinguishes between real and fake images, showcasing our method’s ability to generalize across diverse fake representations. From the visualization, it is evident that NPR struggles with unseen GAN or diffusion models, showing significant overlap between the logits of real and fake categories, often misclassifying fake images as ”real.” In contrast, our method demonstrates superior discrimination, effectively separating ”real” and ”fake” categories, even when encountering unseen sources.
4.5 Ablation Study
To thoroughly evaluate the effectiveness of each component in our method, we conduct ablation experiments by systematically removing some of the proposed modules. The results of these experiments, shown in Table 7, demonstrate the contribution of each module to deepfake detection. Notably, when we remove our main module (LoSFB), the performance of our method drops significantly, falling below 50%. This highlights the crucial role of local spatial-frequency domain feature extraction in detecting forged information.To further validate the effectiveness and rationale behind using the phase component of FFT to extract global information and employing sliding window tiling for local feature extraction in DWT, we remove these modules individually. As shown in the table, the accuracy decreased by 2.4% and 0.8%, respectively. These results emphasize the importance of both effectively extracting local texture information and subtle forged textures embedded in the global information.
5 Conclusion
In this paper, we propose a novel and effective method for deepfake detection, leveraging both local spatial-frequency information and global frequency domain features through two core modules: LoSFB and GloFB. This plug-and-play feature extractor is meticulously designed to capture a broader and more effective set of artifact features. Extensive experiments on 34 diverse generative models strongly demonstrate the generalization capability of our method.
6 Acknowledgments
This work is supported by the National Key R&D Program of China under grant 2021YFB2700900; by the National Natural Science Foundation of China under grant U22B2062, 62172232; by the Jiangsu Basic Research Programs-Natural Science Foundation under grant numbers BK20200039; by the Collaborative Innovation Center of Atmospheric Environment and Equipment Technology (CICAEET) fund.